# Leveraging (Biased) Information: Multi-armed Bandits with Offline Data

### v1

@@ Abstract
We leverage offline data to facilitate online learning in stochastic multi-armed bandits. The probability distributions that govern the offline data and the online rewards can be different. Without any non-trivial upper bound on their difference, we show that no non-anticipatory policy can out-perform the UCB policy by (Auer et al. 2002), even in the presence of offline data. In complement, we propose an online policy MIN-UCB, which outperforms UCB when a non-trivial upper bound is given. MIN-UCB adaptively chooses to utilize the offline data when they are deemed informative, and to ignore them otherwise. MIN-UCB is shown to be tight in terms of both instance independent and dependent regret bounds. Finally, we corroborate the theoretical results with numerical experiments.
Machine Learning, ICML
@@ 1 Introduction
Multi-armed bandit problem (MAB) is a classical model in sequential decision-making. In traditional models, online learning is initialized with no historical dataset. However, in many real world scenarios, historical datasets related to the underlying model exist before online learning begins. Those datasets could be beneficial to online learning. For instance, when a company launch a new product in Singapore, past sales data in the US could be of relevance. This prompts an intriguing question: Can we develop an approach that effectively utilizes offline data, and subsequently learns under bandit feedback? Such approaches carry significant allure, as they potentially reduce the amount of exploration in the learning process. Nevertheless, it is too strong and impractical to assume that the historical dataset and the online model follow the same probability distribution, given the ubiquity of distributional shifts. Therefore, it is desirable to design an adaptive policy that reaps the benefit of offline data if they closely match the online model, while judiciously ignores the offline data and learns from scratch otherwise.
@@
Motivated by the above discussions, we consider a stochastic multi-armed bandit model with possibly biased offline dataset. The entire event horizon consists of a warm-start phase, followed by an online phase whereby decision making occurs. During the warm-start phase, the DM receives an offline dataset, consisting of samples governed by a latent probabiltiy distribution $P^{\text{(off)}}$. The subsequent online phase is the same as standard multi-armed bandit model, except that the offline dataset can be incorporated in decision making. The random online rewards are governed by another latent probabiltiy distribution $P^{\text{(on)}}$. The DM aims to maximize the total cumulative reward earned, or equivalently minimize the regret, during the online phase.
@@
Intuitively, when $P^{\text{(off)}}$, $P^{\text{(on)}}$ are “far apart”, the DM should conduct online learning from scratch and ignore the offline dataset. For example, the vanilla UCB policy by (Auer et al., 2002) incurs expected regret at most
$O\left(\sum_{a:\Delta(a)>0}\frac{\log T}{\Delta(a)}\right),$ (1)
@@
where $T$ is the length of online phase, and $\Delta(a)\geq 0$ is the difference between the expected reward of an optimal arm and arm $a$ (also see Section 2). The bound (1) holds for all $P^{\text{(off)}}$, $P^{\text{(on)}}$. In contrary, when $P^{\text{(off)}}$, $P^{\text{(on)}}$ are “sufficiently close”, the DM should incorporate the offline data into online learning and avoid unnecessary exploration. For example, when $P^{\text{(off)}}=P^{\text{(on)}}$, HUCB1 (Shivaswamy & Joachims, 2012) and MonUCB (Banerjee et al., 2022) incur expected regret at most
$O\left(\sum_{a:\Delta(a)>0}\max\left\{\frac{\log T}{\Delta(a)}-T_{\text{S}}(a)\Delta(a),0\right\}\right),$ (2)
@@
where $T_{\text{S}}(a)$ is the number of offline samples on arm $a$. (2) is a better regret bound than (1). However, (2) only holds when $P^{\text{(off)}}=P^{\text{(on)}}$. The above inspires the following question:
@@
(Q) Can the DM outperform the vanilla UCB, i.e. achieves a better regret bound than (1) when $P^{\text{(off)}}=P^{\text{(on)}}$, while retains the regret bound (1) for general $P^{\text{(off)}},P^{\text{(on)}}$?
@@
The answer to (Q) turns out to be somewhat mixed. Our novel contributions shed light on (Q) as follows:
@@
Impossibility Result. We show in Section 3 that no non-anticipatory policy can achieve the aim in (Q). Even with an offline dataset, no non-anticipatory policy can outperform the vanilla UCB without any additional knowledge or restriction on the difference between $P^{\text{(off)}},P^{\text{(on)}}$.
@@
Algorithm design and analysis. To bypass the impossibility result, we endow the DM with an auxiliary information dubbed valid bias bound $V$, which is an information additional to the offline dataset. The bound $V$ serves as an upper bound on the difference between $P^{\text{(off)}}$ and $P^{\text{(on)}}$. We propose the MIN-UCB policy, which achieves a strictly better regret bound than (1) when $P^{\text{(off)}},P^{\text{(on)}}$ are “sufficiently close”. In particular, our regret bound reduces to (2) if the DM knows $P^{\text{(off)}}=P^{\text{(on)}}$. We provide both instance dependent and independent regret upper bounds on MIN-UCB, and we show their tightness by providing the corresponding regret lower bounds. Our analysis also pins down the precise meaning of “far apart” and “sufficientlly close” in the above discussion. The design and analysis of MIN-UCB are provided in Section 4.
@@
In particular, when $P^{\text{(off)}}=P^{\text{(on)}}$, our analysis provide two novel contributions. Firstly, we establish the tightness of the instance dependent bound (2). Secondly, we establish a pair of instance independent regret upper and lower bounds, which match each other up to a logarithmic factor. MIN-UCB is proved to achieve the upper bound when $P^{\text{(off)}}=P^{\text{(on)}}$. Both the upper and lower bounds are novel in the literature, and we show that the optimal regret bound involves the optimum of a novel linear program.
@@ 1.1 Related Works
Multi-armed bandits with offline data has been studied in multiple works. (Shivaswamy & Joachims, 2012; Banerjee et al., 2022) are the most relevant works, focusing on the special case of $P^{\text{(on)}}=P^{\text{(off)}}$. They only provide instance dependent regret upper bound, while we both instance dependent and independent bounds. Online learning with offline data under $P^{\text{(on)}}=P^{\text{(off)}}$ is also studied in dynamic pricing settings (Bu et al., 2020) and reinforcement learning settings (Hao et al., 2023; Wagenmaker & Pacchiano, 2023), as well as when historical data are sequentially provided (Gur & Momeni, 2022), or provided as clustered data (Bouneffouf et al., 2019; Ye et al., 2020; Tennenholtz et al., 2021). (Zhang et al., 2019) is another closely related work on contextual bandits that allows $P^{\text{(on)}}\neq P^{\text{(off)}}$. We provide more discussions on it in Appendix A.1.
@@
Bayesian policies, such as the Thompson sampling policy (TS), are popular for incorporating offline data into online learning. When the online and offline reward distributions are identical, offline data can be used to construct well-specified priors, which enables better regret bounds than (Russo & Van Roy, 2014; Russo & Roy, 2016; Liu & Li, 2016) when the offline data size increases. By contrast, (Liu & Li, 2016; Simchowitz et al., 2021) show that the TS initialized with mis-specified priors generally incurs worse regret bounds than state-of-the-art online policies without any offline data. Our work focuses on an orthogonal issue of deciding whether to incorporate the possibly biased offline data in online learning.
@@
Our problem model closely relates to machine learning under domain shift. (Si et al., 2023) investigates a offline policy learning problem using historical data on contextual bandits under changing environment from a distributionally robust optimization perspective. (Chen et al., 2022) studies a variant in RL context, and their policy is also endowed with an upper bound on the distributional drift between offline and online models. We provide more discussion in Appendix A.3. Numerous research works consider the issue of distributional drift in supervised learning (Crammer et al., 2008; Mansour et al., 2009; Ben-David et al., 2010) and stochastic optimization (Besbes et al., 2022).
@@
Our work is related to research on online learning with advice, which focuses on improving performance guarantees with hints or predictions. Indeed, our offline dataset could serve as a hint to an online learner. A variety of results are derived in multi-armed bandits (Wei & Luo, 2018), linear bandits (Cutkosky et al., 2022), contextual bandits (Wei et al., 2020b), and full feedback settings (Steinhardt & Liang, 2014; Rakhlin & Sridharan, 2013a, b). These works do not apply in our setting, since they do not consider refining instance dependent bounds. We provide more discussions in Appendix A.2. Our theme of whether to utilize the potentially biased offline data is related to online model selection, studied in (Agarwal et al., 2017; Pacchiano et al., 2020; Lee et al., 2021; Cutkosky et al., 2021; Pacchiano et al., 2023). While these works take a black-box approach, we tailor-make our decision making on whether to utilize offline data with a novel UCB algorithm and the auxilliary input $V$, and the latter is not studied in the abovemoentioned works.

### v2

# Online Bandits with (Biased) Offline Data: Adaptive Learning under Distribution Mismatch

@@ Abstract
Traditional online learning models are typically initialized from scratch. By contrast, contemporary real-world applications often have access to historical datasets that can potentially enhanced the online learning processes. We study how offline data can be leveraged to facilitate online learning in stochastic multi-armed bandits and combinatorial bandits. In our study, the probability distributions that govern the offline data and the online rewards can be different. We first show that, without a non-trivial upper bound on their difference, no non-anticipatory policy can outperform the classical Upper Confidence Bound (UCB) policy by AuerCBF02, even with the access to offline data. In complement, we propose an online policy MIN-UCB for multi-armed bandits. MIN-UCB outperforms the UCB when such an upper bound is available. MIN-UCB adaptively chooses to utilize the offline data when they are deemed informative, and to ignore them otherwise. We establish that MIN-UCB achieves tight regret bounds, in both instance independent and dependent settings. We generalize our approach to the combinatorial bandit setting by introducing MIN-COMB-UCB, and we provide corresponding instance dependent and instance independent regret bounds. We illustrate how various factors, such as the biases and the size of offline datasets, affect the utility of offline data in online learning. We discuss several applications and conduct numerical experiments to validate our findings.
@@ 1  Introduction
Online learning problems with bandit feedback, including multi-armed bandits (MAB), combinatorial bandits and linear bandits, constitute a class of classical models in sequential decision-making. In traditional models, online learning is initialized with no historical dataset. The Decision Maker (DM) must acquire knowledge solely through the online interactions with the environment. By contrast, in many contemporary real-world scenarios, past datasets related to the underlying model are often available before online learning begins. These offline datasets could be potentially beneficial to online learning. For instance, when Apple launches a new iPhone model in Canada, past sales data from the US market could be of relevance. Similar practices are common in operations management, where firms routinely leverage historical demands and pricing data to aid decision-making in new markets or selling seasons.
The above observation raises an intriguing question: Can we develop an algorithm that effectively utilizes offline data for online learning? Such approaches carry significant allure, as they potentially reduce costly exploration in the learning process. Nevertheless, it is often too idealistic, if not flawed, to assume that the historical data and the online environment are governed by the same probability distribution, given the ubiquity of distributional shifts in every-day life applications. Indeed, past data may be biased or misleading. For example, Amazon extrapolated behavioral insights from Kindle Fire tablet users to guide the Fire Phone design. However, the resulting product failed to align with core smartphone demand and was discontinued within a year wohlsen2015amazon. In such cases, using offline data can harm online decision-making.
This presents a fundamental challenge: How to effectively use offline data, that may be biased or mismatch as compared to the online environment? Ideally, we aim to design an adaptive policy that reaps the benefit of offline data when it is informative, i.e. when the offline and online distributions are sufficiently close, while judiciously ignores the offline data and learns from scratch otherwise. Such an approach is particularly valuable in practice, as it can accelerate decision making and reduce costly exploration when offline data is reliable, while avoiding the risks of distribution shifts that have caused failures in real-world applications. Hence, developing methods to manage these two sources of information is both meaningful and necessary.
@@
Motivated by the above discussions, we first consider a stochastic multi-armed bandit model with a possibly biased offline dataset. The learning horizon consists of a warm-start phase, followed by an online phase whereby decision making occurs. During the warm-start phase, the DM receives an offline dataset, consisting of samples governed by a latent probability distribution $P^{\text{(off)}}$. The subsequent online phase is the same as the standard multi-armed bandit model, except that the offline dataset can be incorporated in decision making. The random online rewards are governed by a latent probability distribution $P^{\text{(on)}}$, which can be different from $P^{\text{(off)}}$. The DM aims to maximize the total cumulative reward during the online phase. Subsequently, we extend the above framework to a stochastic combinatorial bandit model.
1.1  Research Intuitions and Main Contributions
@@
Intuitively, in the case of multi-armed bandits, when the offline data distribution $P^{\text{(off)}}$ and online reward distribution $P^{\text{(on)}}$ are “far apart”, the DM should conduct online learning from scratch and ignore the offline dataset. For example, the UCB policy by AuerCBF02, which we call the “vanilla UCB” in the remainder of the manuscript, incurs expected regret (i.e. the performance gap to the optimal policy, formally defined in the forthcoming Section 2) at most
$O\left(\sum_{a\in\mathcal{A}:\Delta(a)>0}\frac{\log T}{\Delta(a)}\right),$ (1)
@@
where $T$ is the number of time steps in the online phase, $\mathcal{A}$ is the arm set, and $\Delta(a)\geq 0$ is the difference between the expected reward of an optimal arm and that of arm $a$ (see Section 2 for the full definition). The bound (1) holds for all $P^{\text{(off)}}$, $P^{\text{(on)}}$. On the contrary, when $P^{\text{(off)}}$ and $P^{\text{(on)}}$ are “sufficiently close”, the DM should incorporate the offline data into online learning and avoid unnecessary exploration. For example, when $P^{\text{(off)}}=P^{\text{(on)}}$, HUCB1 [shivaswamy2012multi] and MonUCB [banerjee2022artificial] incur expected regret at most
$O\left(\sum_{a\in\mathcal{A}:\Delta(a)>0}\max\left\{\frac{\log T}{\Delta(a)}-T_{\text{S}}(a)\Delta(a),0\right\}\right),$ (2)
@@
where $T_{\text{S}}(a)$ is the number of offline samples on arm $a$. (2) is a strictly smaller regret bound than (1). However, (2) only holds when $P^{\text{(off)}}=P^{\text{(on)}}$. The above inspires the following question:
@@
(Q) Can the DM outperform the vanilla UCB, i.e. achieves a strictly smaller regret bound than (1) when $P^{\text{(off)}}=P^{\text{(on)}}$, while retains the regret bound (1) for general $P^{\text{(off)}},P^{\text{(on)}}$?
@@
The answer to (Q) turns out to be somewhat mixed. Our novel contributions shed light on (Q):
@@
An Impossibility Result. In Section 3, we show that no non-anticipatory policy can achieve the aim in (Q). Even with an offline dataset, no non-anticipatory policy can outperform the vanilla UCB without any additional knowledge or restriction on the difference between $P^{\text{(off)}},P^{\text{(on)}}$.
@@
Multi-armed Bandit Models. To bypass the impossibility result, we assume that the DM is provided with auxiliary information dubbed valid bias bound $V$, which supplements the offline dataset. The bound $V$ serves as an upper bound on the difference between $P^{\text{(off)}}$ and $P^{\text{(on)}}$. We propose the MIN-UCB policy, which achieves a strictly smaller regret bound than (1) when $P^{\text{(off)}},P^{\text{(on)}}$ are “sufficiently close”. In particular, our regret bound reduces to (2) if the DM knows $P^{\text{(off)}}=P^{\text{(on)}}$. We provide both instance-dependent and instance-independent regret upper bounds on MIN-UCB. Additionally, we show the tightness of both regret upper bounds by providing the corresponding regret lower bounds. Our analysis precisely characterizes the meaning of “far apart” and “sufficiently close” in the above discussion. The design and analysis of MIN-UCB are provided in Section 4.
Generalization to Combinatorial Bandit Models. We extend our framework to combinatorial bandits in Section 5. We propose the MIN-COMB-UCB policy, and establish an instance-independent regret upper bound. Similar to the multi-armed bandit case, our regret upper bound is uniformly better than that without access to offline data, such as chen2013combinatorial. When $P^{\text{(off)}},P^{\text{(on)}}$ are sufficiently close, our regret upper bound is strictly smaller than that in chen2013combinatorial. In the special case of combinatorial bandits with linear rewards, we further derive tight instance-independent regret upper bounds and instance-dependent regret upper bounds. The latter shares a similar saving term to that in the multi-armed bandit model, quantifying the benefit of offline data. In particular, when we specialize the combinatorial bandit model degenerates to the multi-armed bandit model, our regret upper bounds reduce to those presented in Section 4. The generalization shows the adaptivity and scalability of our approach to broader online learning models.
Technical Novelty. Our technical contributions are summarized as follows.
  1. 1.
    Impossibility Result under Unknown Bias: To the best of our knowledge, our work is the first to establish a regret lower bound in the presence of potentially biased offline data, without knowing any prior information on the discrepancy between $P^{\text{(off)}}$ and $P^{\text{(on)}}$.
  2. 2.
@@
    Tight Instance-dependent Regret Bounds: When $P^{\text{(off)}}=P^{\text{(on)}}$, we establish the tightness of the instance-dependent regret upper bound (2), which is novel in the literature. We further provide matching instance-dependent upper and lower bounds for general $P^{\text{(off)}},P^{\text{(on)}}$, precisely quantifying the impact of $T_{\text{S}}(a)$ and $V$ on each sub-optimal arm $a$. The tight analysis also represents new contributions to the literature.
  3. 3.
    Tight Instance-independent Regret Bounds: We establish a pair of instance-independent regret upper and lower bounds, which match each other up to a logarithmic factor. This optimal regret bound quantifies the effect of arbitrary $\{T_{\text{S}}(a)\}_{a\in\mathcal{A}}$ and its heterogeneity by involving the optimum of a novel linear program. Even in the special case when $P^{\text{(off)}}=P^{\text{(on)}}$, both the upper and lower bounds are novel in the literature.
Finally, we provide numerical experiments in Section 6.
1.2  Key Insights and Managerial Implications
Our technical contributions bear several important implications. Firstly, our study sheds light on the benefits of offline data in multi-armed bandit settings, where the offline data are potentially distributed differently from the online data. The takeaway message is: the DM can only reap the benefits of offline data, when the DM has non-trivial auxiliary information about the discrepancy between offline and online reward distributions. The necessity of auxiliary information is formalized in our impossibility result in Section 3. The following thought experiment provides another perspective. Consider multi-armed bandits with possibly biased offline data, but without any auxiliary information. Now, it is intuitive to conduct an exploration to identify whether the offline data helps us accelerate the estimation of the mean rewards. Our takeaway message is that the cost of such an exploration outweighs the benefits of the offline data. By contrast, by harnessing suitable auxiliary information, we can tip the balance and reap the benefits of offline data.
Secondly, offline data improve the performance over the vanilla UCB only when the reward distributions $P^{\text{(off)}}$ and $P^{\text{(on)}}$ are sufficiently close. A main contribution of our analysis is the identification of a discrepancy measure between the offline and online distributions for each arm $a$. When this discrepancy is large, the offline data should be disregarded; when it is small, the offline data becomes informative and can be effectively utilized to enhance performance. In the latter case, MIN-UCB achieves a strictly smaller regret bound than the vanilla UCB. The improvement, captured by a saving term, becomes larger when $T_{\text{S}}(a)$ for an $a\in{\cal A}$ increases, and $V(a)$ for an $a\in{\cal A}$ decreases. A similar phenomenon appears in our instance-independent regret bound. Only when $V(a)$’s are sufficiently small can the minimax regret bound be improved. Moreover, the improvement becomes more pronounced when $\min_{a\in\mathcal{A}}T_{\text{S}}(a)$ is larger and $V(a)$ for each $a$ is smaller, reflecting the intuition that a larger and more reliable offline dataset enables more effective online learning.
Thirdly, when the offline data is informative, factors other than the size of offline dataset can influence algorithm performance. In particular, the performance depends not only on the value of $V$ and discrepancy between $P^{\text{(off)}}$ and $P^{\text{(on)}}$, but also on the direction of the discrepancy. For example, if $P^{\text{(off)}}$ underestimates the expected reward of a sub-optimal arm $a$, it allows our algorithm to eliminate $a$ faster, even if $P^{\text{(off)}}$ and $P^{\text{(on)}}$ are “far apart”. Besides, the heterogeneity in offline dataset sizes $\{T_{\text{S}}(a)\}_{a\in\mathcal{A}}$ also plays a critical role. The impact is not determined solely by the minimum, maximum, or average of $\{T_{\text{S}}(a)\}_{a\in\mathcal{A}}$, but rather depends on the full profile of these values in a non-trivial way.
Lastly, when the offline data is informative, our approach yields notable improvements in several important special cases. Specifically, consider the setting where abundant offline data is available for only a subset of arms, while no prior information exists for the remaining arms. In other words, the DM has reliable information on a small set but must make decisions across a much larger action space. Another case arises when the length of online phase $T$ is relatively small compared to the number of arms $K$ and $\{T_{\text{S}}(a)\}_{a\in\mathcal{A}}$, such that the exploration budget is severely limited. In both scenarios, our policy demonstrates explicit and significant improvements over the standard online learning algorithms without access to the offline data. See Section 4 for more details.
@@ 1.3  Related Works
Online Learning with Offline Data The study of multi-armed bandits with offline data has attracted increasing attention. shivaswamy2012multi, banerjee2022artificial are the most relevant works, analyzing the special case of $P^{\text{(on)}}=P^{\text{(off)}}$ and establishing only the instance dependent regret upper bound, whereas we provide both tight instance dependent and independent regret bounds. Online learning with offline data under $P^{\text{(on)}}=P^{\text{(off)}}$ is also explored in dynamic pricing bu2020online and reinforcement learning HaiJLVW23, WagenmakerP23. Extensions include sequentially revealed historical data gur2022adaptive and clustered data bouneffouf2019optimal, ye2020combining, tennenholtz2021bandits. zhang2019warm is another closely related work on contextual bandits that allows $P^{\text{(on)}}\neq P^{\text{(off)}}$. We provide a technical discussion in Appendix A.1.
@@
Bayesian Policies Thompson sampling (TS) and related Bayesian policies are widely used to integrate offline data into online learning. When the online and offline reward distributions are identical, offline data can be used to construct well-specified priors that yield improved regret bounds than russo2014learning, RussoV16, LiuL16 when the offline data size increases. In contrast, [LiuL16, SimchowitzTKHLDS21] show that TS with misspecified priors may perform worse than state-of-the-art online policies without offline data. Our paper addresses the orthogonal issue of determining whether to incorporate possibly biased offline data in online learning.
@@
Machine Learning under Domain Shift Our model is closely related to machine learning under domain shift. SiNZB23 considers offline policy learning with historical data on contextual bandits setting in a changing environment from a distributionally robust optimization perspective. chen2022data explores a variant in RL context, whose policy is also endowed with an upper bound on the distributional drift between offline and online models. Further discussion is provided in Appendix A.3. Generally, distributional drift has been extensively studied in supervised learning CrammerKW08, mansour2009domain, BenDavidBCKPV10 and stochastic optimization besbesMM22.
@@
Online Learning with Advice Our work is related to online learning with advice, which considers improving performance guarantees with hints or predictions. The offline dataset in our setting can be viewed as such a hint. Prior works include multi-armed bandits WeiL18, linear bandits CutkoskyDDZ22, contextual bandits wei2020taking, full feedback settings steinhardt2014adaptivity, rakhlin2013online, rakhlin2013optimization, online episodic MDPs golowich2022can, and discounted MDPs feng2019does. These works do not apply to our setting since they do not refine instance dependent regret bounds. We provide more discussions in Appendix A.2. Our theme of whether to leverage potentially biased offline data is also related to online model selection AgarwalLNS17, pacchiano2020model, lee2021online, cutkosky2021dynamic, PacchianoDG23. Unlike these black-box approaches, we tailor our decision-making on whether to use offline data with a novel UCB algorithm and auxiliary input $V$, and the latter is not studied in the above-mentioned works.
Comparison to the Conference Version cheung2024leveraging A preliminary version of this manuscript appears in cheung2024leveraging. Our work incorporates the following additional contributions compared to cheung2024leveraging:
  • Extentions to Combinatorial Bandits. cheung2024leveraging only studies the multi-armed bandit setting. In the current manuscript, we extend our framework to general combinatorial bandits. We propose a new policy, MIN-COMB-UCB, which allows efficient (computationally and regret-wise) online learning even when the number of combinatorial arms can be exponential in the number of basic arms. For MIN-COMB-UCB, we establish an instance-independent regret upper bound, which is strictly smaller than the existing bounds without offline data when $P^{(\text{off})}$, $P^{(\text{on})}$ are close enoguh. For the special case with linear rewards case, we provide tight instance-independent regret bounds and an instance-dependent regret upper bound with a similar “Saving” term. All these regret bounds rely on novel techniques that combine the standard combinatorial bandit framework with an arbitrary offline dataset.
  • Discussions on Applications. We discuss applications of our framework in the context of dynamic pricing and social influence maximization, illustrating how our framework can be applied to these real-world scenarios. We demonstrate that leveraging offline data alongside online learning is particularly valuable for improving decision quality and scalability.
  • Numerical Experiments. We conduct more systematic experiments to examine the effect of discrepancy, the length of online phase $T$, and the size of offline dataset on algorithm performance. The results validate our theoretical findings and show that the bias, direction, and magnitude of the offline dataset all affect performance. Furthermore, the results highlights the importance of both high-quality data (small bias) and effective algorithms (such as our MIN-UCB).
1.4  Notation